# Estimation of discrete distributions in relative entropy, and the deviations of the missing mass

### v2

@@ 6.1 Risk decomposition
Using that, for any $u\in\mathbb{R}^{+}$, one has $31-(1+u)^{-1/2}=u/(1+u+\sqrt{1+u})\leqslant u/(2\sqrt{u}+\sqrt{u})=\sqrt{u}/$ 3, we can bound the first term in (61) as
@@ 6.2 Upper bound in Hellinger distance
In this section, we proceed with the control of the first term in the decomposition of Lemma 3. Specifically, Lemma 4 below provides a high-probability bound on the squared Hellinger distance between the empirical distribution $\overline{P}_{n}$ and the true distribution $P$. This bound is established by combining the the analysis of the reverse-relative entropy from [Agr22] (for large probabilities) with concentration inequalities deduced from negative association [JDP83] (for small probabilities).
@@
We now turn to the second term in (65). Note that $\sum_{j\in J^{-}}\overline{p}_{j}=n^{-1}\sum_{j\in J^{-}}N_{j}$. In addition, by [JDP83, 3.1(a)], the variables $(N_{j})_{1\leqslant j\leqslant d}$ are negatively associated, in the sense of [JDP83, Definition 2.1]. This implies in particular that, for every $t\in\mathbb{R}^{+}$,
$\mathbb{E}\bigg{[}\exp\bigg{(}t\sum_{j\in J^{-}}N_{j}\bigg{)}\bigg{]}\leqslant\prod_{j\in J^{-}}\mathbb{E}\big{[}e^{tN_{j}}\big{]}=\prod_{j\in J^{-}}\big{(}1-p_{j}+p_{j}e^{t}\big{)}^{n}$
$\leqslant\prod_{j\in J^{-}}\exp\big{\{}np_{j}(e^{t}-1)\big{\}}=\exp\bigg{\{}n\sum_{j\in J^{-}}p_{j}(e^{t}-1)\bigg{\}}=\exp\big{\{}s^{-}(e^{t}-1)\big{\}}\,.$
Rearranging and using that $e^{t}-1-t=\sum_{k\geqslant 2}\frac{t^{k}}{k!}\leqslant\sum_{k\geqslant 2}\frac{t^{k}}{2}=\frac{t^{2}}{2(1-t)}$ for any $t\in[0,1)$, we obtain that for any $10\leqslant t<$ 1:
$\log\mathbb{E}\bigg{[}\exp\bigg{(}t\,\bigg{\{}\sum_{j\in J^{-}}N_{j}-s^{-}\bigg{\}}\bigg{)}\bigg{]}\leqslant s^{-}(e^{t}-1-t)\leqslant\frac{s^{-}t^{2}}{2(1-t)}\,.$
Using again the sub-gamma tail bound [BLM13, p. 29], we conclude that, with probability at least $1-\delta$,
$\sum_{j\in J^{-}}N_{j}<s^{-}+\sqrt{2s^{-}\log(1/\delta)}+\log(1/\delta)\leqslant 2\big{(}s^{-}+\log(1/\delta)\big{)}\,,$
namely
$\sum_{j\in J^{-}}\big{(}p_{j}+\overline{p}_{j}\big{)}=n^{-1}\bigg{(}s^{-}+\sum_{j\in J^{-}}N_{j}\bigg{)}<\frac{3s^{-}+2\log(1/\delta)}{n}\,.$ (68)
@@ 6.3 Control of the contribution of underestimated frequencies
Roughly speaking, inequality (69) will be used in the analysis of confidence-independent estimators such as the Laplace estimator (or of confidence-dependent estimators in the low confidence regime), while (70) will be used in the analysis of the confidence-dependent estimators in the high-confidence regime.
@@
We now aim to obtain a high-probability bound on the sum $\sum_{j=1}^{d_{\lambda}}W_{j}^{(\lambda)}$, which is a sum of i.i.d. nonnegative random variables. Unfortunately, since $W^{(\lambda)}$ has super-exponential tails, its moment generating functions is infinite at any positive value, thus one cannot rely on the Chernoff approach to obtain such a tail bound. We will instead work with its moments, using a moment estimate of Latała [Lat97]. Specifically, Lemma 8 below is a consequence of (part of) [Lat97, Corollary 1] obtained by tracking the numerical constants in this result. We recall that, for any $p\in\mathbb{R}^{+}$ such that $p\geqslant 1$ and any real random variable $Z$, we let $\|Z\|_{p}=\mathbb{E}[|Z|^{p}]^{1/p}\in\mathbb{R}^{+}\cup\{+\infty\}$.
@@
we deduce that $\phi^{\prime}_{\lambda}(t)<0$ for any $t\geqslant 50p$, thus $\phi_{\lambda}$ decreases over $[\max\{50p,4\lambda\},+\infty)$.
@@ 6.6 Proof of Theorem 5
$R_{\widetilde{\lambda}}=\sum_{j\,:\;p_{j}\geqslant 4\widetilde{\lambda}/n}p_{j}\log({2np_{j}})\bm{1}\Big{(}N_{j}\leqslant\frac{np_{j}}{4}\Big{)}+\sum_{j\,:\;p_{j}\geqslant 4\widetilde{\lambda}/n}p_{j}\bm{1}\Big{(}N_{j}\leqslant\frac{np_{j}}{4}\Big{)}\cdot\log\bigg{(}\frac{1}{\widetilde{\lambda}}\bigg{)}\,.$ (87)
@@
[…]
@@ 7.2 Proof of Lemma 1 and Theorem 4
=$\mathcal{F}=\Big{\{}P^{(j)}=\delta^{1/n}\delta_{1}+\big{(}1-\delta^{1/n}\big{)}\delta_{j}:1\leqslant j\leqslant d\Big{\}}=\big{\{}\delta_{1}\big{\}}\cup\big{\{}P^{(j)}:2\leqslant j\leqslant d\big{\}}\,.$
@@ 7.3 Proof of Proposition 1
Note that if $s\leqslant 35$, then $01-3e^{-s/35}<$ 0 and the inequality is trivial. From now on, we therefore assume that $s\geqslant 36$.
@@
(Indeed, for any $\sigma\in\mathcal{S}\setminus\widehat{\mathcal{S}}$, the distribution $P_{\sigma}^{\otimes n}$ puts a mass of 00 to the sequence $(X_{1},\dots,X_{n})$; while all measures $P_{\sigma}^{\otimes n}$ with $\sigma\in\widehat{\mathcal{S}}$ put the same positive mass to such a sequence.) In other words, $\sigma=\widehat{\sigma}\cup\widetilde{\sigma}$, where (conditionally on $X_{1},\dots,X_{n}$) $\widetilde{\sigma}$ is uniformly distributed over all subsets of $\{2,\dots,d\}\setminus\widehat{\sigma}$ with $s-D_{n}$ elements. Write $\widehat{P}_{n}=(\widehat{p}_{1},\dots,\widehat{p}_{d})$, and let $\widehat{\alpha}=n(1-\widehat{p}_{1})$.
@@
$\mathrm{KL}(P_{\sigma},\widehat{P}_{n})\geqslant D\bigg{(}\sum_{j=2}^{d}p_{j}^{\sigma},\sum_{j=2}^{d}\widehat{p}_{j}\bigg{)}=D\bigg{(}\frac{s-1}{2en},\frac{\widehat{\alpha}}{n}\bigg{)}=\frac{\widehat{\alpha}}{n}\cdot h\Big{(}\frac{s-1}{2e\widehat{\alpha}}\Big{)}\,.$
@@
$\mathrm{KL}(P_{\sigma},\widehat{P}_{n})$ $=\sum_{j=1}^{d}D(p_{j}^{\sigma},\widehat{p}_{j})\geqslant\sum_{2\leqslant j\leqslant d,\,j\not\in\widehat{\sigma}}D\big{(}p_{j}^{\sigma},\widehat{p}_{j}\big{)}\geqslant\sum_{2\leqslant j\leqslant d,\,j\not\in\widehat{\sigma}}D\Big{(}\frac{1}{2en},\widehat{p}_{j}\Big{)}\bm{1}(j\in\widetilde{\sigma})\,.$ (95)
@@
$|\widehat{J}|$ $\geqslant(d-1)-\frac{d}{2}-|\widehat{\sigma}|\geqslant\frac{d}{2}-1-(e-1)-\frac{s-1}{2}=\frac{d-s-(2e-1)}{2}$
@@
$\mathrm{KL}(P_{\sigma},\widehat{P}_{n})\geqslant\frac{1}{4e^{2}n}\log\bigg{(}\frac{ed}{s}\bigg{)}\sum_{j\in\widehat{J}}\bm{1}(j\in\widetilde{\sigma})=\frac{1}{4e^{2}n}\log\bigg{(}\frac{ed}{s}\bigg{)}|\widehat{J}\cap\widetilde{\sigma}|\,.$ (97)
@@ 8 Proof of Theorem 6
[…]
Now since $\mathbb{P}(\widetilde{N}_{j}\leqslant\lambda_{j}/4)\leqslant e^{-\lambda_{j}/14}$, the (Bernoulli) random variable $\bm{1}(\widetilde{N}_{j}\leqslant\lambda_{j}/4)$ is stochastically dominated by $P_{j}^{\prime}\sim\mathsf{Poisson}(e^{-\lambda_{j}/14})$. By choosing $(P_{j}^{\prime})_{1\leqslant j\leqslant d}$ to be independent, we deduce from Lemma 18 that, for every $k\in\mathbb{N}$,
[…]
@@
Since $P_{k}\sim\mathsf{Poisson}(d_{k}^{\prime}e^{-2^{k}/14})$, the Poisson tail bound (Lemma 17) implies that for any $u\geqslant d_{k}^{\prime}e^{-2^{k}/28}$,
$\mathbb{P}\big{(}P_{k}\geqslant eu\big{)}\leqslant\exp\big{\{}-D\big{(}eu,d_{k}^{\prime}e^{-2^{k}/14}\big{)}\big{\}}\,.$
Now since $eu/(d_{k}^{\prime}e^{-2^{k}/14})\geqslant e(d_{k}^{\prime}e^{-2^{k}/28})/(d_{k}^{\prime}e^{-2^{k}/14})\geqslant e\cdot e^{2^{k}/28}$, and since $D(p,q)\geqslant e^{-1}p\log(p/q)$ when $p\geqslant eq$ (Lemma 15), the previous bound implies that
@@ Conclusion.
[…]
@@ 9 Proof of Proposition 2
Let $X_{n+1}$ be a new observation from $P$, independent from $X_{1},\dots,X_{n}$. For $j=1,\dots,d$, denote by $\widetilde{N}_{j}=\sum_{i=1}^{n+1}\bm{1}(X_{i}=j)$ the count of class $j$ and by $\widetilde{D}=\sum_{j=1}^{d}\bm{1}(\widetilde{N}_{j}\geqslant 1)$ the number of distinct classes in the extended sample $(X_{1},\dots,X_{n+1})$. Also, let $\widetilde{N}=(\widetilde{N}_{j})_{1\leqslant j\leqslant d}$. Since the joint distribution of $(X_{1},\dots,X_{n+1})$ is exchangeable (and permuting indices does not change $\widetilde{N}$), for every $j=1,\dots,d$ and $i=1,\dots,n$ one has $\mathbb{P}(X_{n+1}=j|\widetilde{N})=\mathbb{P}(X_{i}=j|\widetilde{N})$; hence,
$\mathbb{P}(X_{n+1}=j|\widetilde{N})$ $=\frac{1}{n+1}\sum_{i=1}^{n}\mathbb{P}(X_{i}=j|\widetilde{N})=\frac{1}{n+1}\mathbb{E}\bigg{[}\sum_{i=1}^{n+1}\bm{1}(X_{i}=j)\Big{|}\widetilde{N}\bigg{]}$
@@
$=\log\bigg{(}\frac{\widetilde{N}_{j}}{\widetilde{N}_{j}-1}\bigg{)}\cdot\mathbb{P}(X_{n+1}=j|\widetilde{N})$
@@
and in addition $\mathbb{E}_{P}[\widetilde{C}_{2}]\leqslant\mathbb{E}_{P}[D_{n}]\leqslant s_{n}=s_{n}(P)$. Thus
@@
Plugging this inequality into (119), and using that the map $x\mapsto x\log(ed/x)$ is increasing over $[0,d]$, leads to the desired bound (32).
@@ 10 Technical lemmata
We conclude with the proof of the bound $h(t)\leqslant(t-1)^{2}$. To this end, let $f(t)=h(t)/(t-1)^{2}$ for $t\in\mathbb{R}^{+}\setminus\{1\}$, and $f(1)=1/2$. It is easily verified that $f$ is continuous on $\mathbb{R}^{+}$ and differentiable on +$\mathbb{R}^{+}\setminus\{1\}$. In addition, $f^{\prime}(t)=g(t)/(t-1)^{3}$ where $g(t)=2(t-1)-(t+1)\log t$ for $t>0$. One has $g^{\prime}(t)=-(t^{-1}-1-\log(t^{-1}))\leqslant 0$, and since $g(1)=0$, this implies that $g\leqslant 0$ on $(0,1]$ while $g\geqslant 0$ on $[1,+\infty)$. Therefore $f^{\prime}(t)=g(t)/(t-1)^{3}\leqslant 0$ for any $t\in\mathbb{R}^{+}\setminus\{1\}$ and thus $f$ is decreasing on $\mathbb{R}^{+}$. We conclude by noting that $\lim_{t\to 0^{+}}f(t)=1$, so that $f\leqslant 1$ on $\mathbb{R}_{+}^{*}$. ∎
@@
For any $p\in\mathbb{R}_{+}^{*}$, the function $q\mapsto D(p,q)$ is strictly convex on $\mathbb{R}_{+}^{*}$, and reaches its minimum (equal to 00) at $q=p$. Hence, it is decreasing on $(0,p]$ and increasing on $[p,+\infty)$. In addition, $D(p,q)\leqslant p\log(p/q)$ when $q\leqslant p$, and $D(p,q)\geqslant e^{-1}p\log(p/q)$ when $q\leqslant p/e$.
@@
We may assume, without loss of generality, that $q_{j}>0$ for any $j\in J$. Indeed, either there exists a $j\in J$ such that $q_{j}=0$ and $p_{j}>0$, in which case the left-hand side of (120) is +$\infty$+ ∞ and the inequality holds; or, for any $j\in J$ such that $q_{j}=0$, one has $p_{j}=0$: but in this case replacing $J$ by $J^{\prime}=\{j\in J:q_{j}>0\}$ does not affect the right-hand side of (120).
Next, if $J=\varnothing$, the right-hand side of (120) is 00, thus the inequality holds. We now also assume that $J\neq\varnothing$. Then, by non-negativity and convexity of the function $h$ (Lemma 14), we have
@@
If $q=0$, then by assumption $p=0$ as well, hence all terms in (126) are equal to 00 and the inequalities hold; the same holds when $p=q$. We thus assume $q>0$ and $p\neq q$. In this case, we have $D(p,q)/(\sqrt{p}-\sqrt{q})^{2}=\phi(p/q)$. Now, a direct computation shows that the map $t\mapsto\phi(t^{2})$ is continuously differentiable on $\mathbb{R}_{+}^{*}$, with derivative $t\mapsto 2(t-1)^{-3}\psi(t)$ where $\psi(t)=t^{2}-2t\log t-1$. Since $\psi^{\prime}(t)=2(t-1-\log t)\geqslant 0$ for any $t>0$, the function $\psi$ is non-decreasing, and since $\psi(1)=0$ we deduce that $\psi\leqslant 0$ on $(0,1]$ and $\psi\geqslant 0$ on $[1,+\infty)$. Hence $\frac{\mathrm{d}}{\mathrm{d}t}\phi(t^{2})=2(t-1)^{-3}\psi(t)\geqslant 0$ for any $t>0$, $t\neq 1$, thus by continuity of $\phi$ at 00 and $1$, the function $\phi$ is non-decreasing on $\mathbb{R}^{+}$.

### v3

@@ 6.1 Risk decomposition
Using that, for any $u\in\mathbb{R}^{+}$, one has $1-(1+u)^{-1/2}=u/(1+u+\sqrt{1+u})\leqslant u/(2\sqrt{u}+\sqrt{u})=\sqrt{u}/3$, we can bound the first term in (61) as
@@ 6.2 Upper bound in Hellinger distance
In this section, we proceed with the control of the first term in the decomposition of Lemma 3. Specifically, Lemma 4 below provides a high-probability bound on the squared Hellinger distance between the empirical distribution $\overline{P}_{n}$ and the true distribution $P$. This bound follows from the analysis of the reverse-relative entropy from [Agr22] (for large probabilities) combined with a simple binomial deviation bound (for small probabilities).
@@
We now turn to the second term in (65). Note that
$n\sum_{j\in J^{-}}\overline{p}_{j}=\sum_{j\in J^{-}}N_{j}=\sum_{i=1}^{n}\sum_{j\in J^{-}}\bm{1}(X_{i}=j)=\sum_{i=1}^{n}\bm{1}(X_{i}\in J^{-})\,$,
which follows a binomial distribution with parameters $n$ and $\sum_{j\in J^{-}}p_{j}=s^{-}/n$. Bernstein’s inequality [BLM13, Theorem 2.10 p. 37] then yields, with probability at least $1-\delta$,
$\sum_{j\in J^{-}}\big(p_{j}+\overline{p}_{j}\big)<\frac{s^{-}}{n}+\frac{s^{-}+\sqrt{2s^{-}\log(1/\delta)}+\log(1/\delta)}{n}<\frac{3s^{-}+2\log(1/\delta)}{n}\,.$ (68)
@@ 6.3 Control of the contribution of underestimated frequencies
Roughly speaking, inequality (69) will be used in the analysis of confidence-independent estimators such as the Laplace estimator (or of confidence-dependent estimators in the low confidence regime), while (70) will be used in the analysis of the confidence-dependent estimators in the high-confidence regime. Intuitively, the first term in (69) comes from the contribution of all classes, while the second term may come from single class $j$ with $p_{j}\asymp\log(1/\delta)/n$.
@@
We now aim to obtain a high-probability bound on $\sum_{j=1}^{d_{\lambda}}W_{j}^{(\lambda)}$, which is a sum of i.i.d. nonnegative random variables. Unfortunately, since $W^{(\lambda)}$ has super-exponential tails, its moment generating functions is infinite at any positive value, thus one cannot rely on the Chernoff approach to obtain such a tail bound. We will instead work with its moments, using a moment estimate of Latała [Lat97]. Specifically, Lemma 8 below is a consequence of (part of) [Lat97, Corollary 1] obtained by tracking the numerical constants in this result. We recall that, for any $p\in\mathbb{R}^{+}$ such that $p\geqslant 1$ and any real random variable $Z$, we let $\|Z\|_{p}=\mathbb{E}[|Z|^{p}]^{1/p}\in\mathbb{R}^{+}\cup\{+\infty\}$.
@@
we deduce that $\phi^{\prime}_{\lambda}(t)<0$ for any $t\geqslant 50p$, thus $\phi_{\lambda}$ decreases over $\max\{50p,4\lambda\},+\infty)$.
@@ 6.6 Proof of Theorem 5
[…]
@@
$=\sum_{j\,:\;4\widetilde{\lambda}/n\leqslant p_{j}<4/n}p_{j}\log({2np_{j}})\bm{1}\Big(N_{j}\leqslant\frac{np_{j}}{4}\Big)+\sum_{j\,:\;p_{j}\geqslant 4/n}p_{j}\log({2np_{j}})\bm{1}\Big(N_{j}\leqslant\frac{np_{j}}{4}\Big)$
$\leqslant\log(8)\sum_{j\,:\;4\widetilde{\lambda}/n\leqslant p_{j}<4/n}p_{j}\bm{1}\Big(N_{j}\leqslant\frac{np_{j}}{4}\Big)+R_{1}\leqslant\log(8)\sum_{j\,:\;p_{j}<4/n}p_{j}+R_{1}$
@@ 7.2 Proof of Lemma 1 and Theorem 4
$\mathcal{F}=\Big\{P^{(j)}=\delta^{1/n}\delta_{1}+\big(1-\delta^{1/n}\big)\delta_{j}:1\leqslant j\leqslant d\Big\}=\big\{\delta_{1}\big\}\cup\big\{P^{(j)}:2\leqslant j\leqslant d\big\}\,.$
@@ 7.3 Proof of Proposition 1
Note that if $s\leqslant 35$, then $1-3e^{-s/35}<0$ and the inequality is trivial. From now on, we therefore assume that $s\geqslant 36$.
@@
(Indeed, for any $\sigma\in\mathcal{S}\setminus\widehat{\mathcal{S}}$, the distribution $P_{\sigma}^{\otimes n}$ puts a mass of 0 to the sequence $(X_{1},\dots,X_{n})$; while all measures $P_{\sigma}^{\otimes n}$ with $\sigma\in\widehat{\mathcal{S}}$ put the same positive mass to such a sequence.) In other words, $\sigma=\widehat{\sigma}\cup\widetilde{\sigma}$, where (conditionally on $X_{1},\dots,X_{n}$) $\widetilde{\sigma}$ is uniformly distributed over all subsets of $\{2,\dots,d\}\setminus\widehat{\sigma}$ with $s-D_{n}$ elements. Write $\widehat{P}_{n}=(\widehat{p}_{1},\dots,\widehat{p}_{d})$, and let $\widehat{\alpha}=n(1-\widehat{p}_{1})$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 8 Proof of Theorem 6
$\sum_{j\,:\;\lambda_{j}\geqslant 1}\lambda_{j}\bm{1}\Big(\widetilde{N}_{j}\leqslant\frac{\lambda_{j}}{4}\Big)$ $=\sum_{k\geqslant 0}\sum_{j\,:\;2^{k}\leqslant\lambda_{j}<2^{k+1}}\lambda_{j}\bm{1}\Big(\widetilde{N}_{j}\leqslant\frac{\lambda_{j}}{4}\Big)$
$\leqslant\sum_{k\geqslant 0}2^{k+1}\sum_{j\,:\;2^{k}\leqslant\lambda_{j}<2^{k+1}}\bm{1}\Big(\widetilde{N}_{j}\leqslant\frac{\lambda_{j}}{4}\Big)\,.$
Now, let $d_{k}^{\prime}=|\{1\leqslant j\leqslant d:2^{k}\leqslant\lambda_{j}<2^{k+1}\}|$ for each $k\geqslant 0$, and let $(P_{k})_{k\geqslant 0}$ be independent random variables with $P_{k}$ following the binomial distribution $\mathsf{Bin}(d_{k}^{\prime},e^{-2^{k}/14})$. Since $\mathbb{P}(\widetilde{N}_{j}\leqslant\lambda_{j}/4)\leqslant e^{-\lambda_{j}/14}\leqslant e^{-2^{k}/14}$ for each $k\geqslant 0$ and $j$ such that $2^{k}\leqslant\lambda_{j}<2^{k+1}$, and since the random variables $\widetilde{N}_{j}$ are independent, it follows from Lemma 18 that
$\sum_{j\,:\;2^{k}\leqslant\lambda_{j}<2^{k+1}}\bm{1}\Big(\widetilde{N}_{j}\leqslant\frac{\lambda_{j}}{4}\Big)\preccurlyeq P_{k}\,.$
The previous inequalities together with another application of Lemma 18 imply that
@@
Since $P_{k}\sim\mathsf{Bin}(d_{k}^{\prime},e^{-2^{k}/14})$, Bennett’s inequality [BLM13, Theorem 2.9 p. 35] implies that for any $u\geqslant d_{k}^{\prime}e^{-2^{k}/28}$,
$\mathbb{P}\big(P_{k}\geqslant eu\big)\leqslant\exp\bigg\{-d_{k}^{\prime}e^{-2^{k}/14}h\bigg(\frac{eu}{d_{k}^{\prime}e^{-2^{k}/14}}\bigg)\bigg\}\,.$
Now since $eu/(d_{k}^{\prime}e^{-2^{k}/14})\geqslant e(d_{k}^{\prime}e^{-2^{k}/28})/(d_{k}^{\prime}e^{-2^{k}/14})\geqslant e\cdot e^{2^{k}/28}$, and since $h(t)\geqslant e^{-1}t\log t$ when $t\geqslant e$ (Lemma 14), the previous bound implies that
@@ Conclusion.
$\sum_{k\geqslant 0}d_{k}^{\prime}e^{-2^{k}/56}=\sum_{k\geqslant 0}\sum_{j\,:\;2^{k}\leqslant np_{j}<2^{k+1}}e^{-2^{k}/56}$
$\leqslant\sum_{k\geqslant 0}\sum_{j\,:\;2^{k}\leqslant np_{j}<2^{k+1}}e^{-np_{j}/112}=\sum_{j\,:\;p_{j}\geqslant 1/n}e^{-np_{j}/112}\,.$
@@ 9 Proof of Proposition 2
[…]
@@
[…]
@@
and in addition $\mathbb{E}_{P}[\widetilde{C}_{2}]\leqslant\mathbb{E}_{P}[D_{n}]\leqslant s_{n}=s_{n}(P)$. Using that the map $x\mapsto x\log(ed/x)$ is increasing over $[0,d]$, we deduce that
@@
Plugging this inequality into (119) and bounding $1+\log 4\leqslant 2.4$ leads to the desired bound (32).
@@ 10 Technical lemmata
We conclude with the proof of the bound $h(t)\leqslant(t-1)^{2}$. To this end, let $f(t)=h(t)/(t-1)^{2}$ for $t\in\mathbb{R}^{+}\setminus\{1\}$, and $f(1)=1/2$. It is easily verified that $f$ is continuous on $\mathbb{R}^{+}$ and differentiable on $\mathbb{R}^{+}\setminus\{1\}$. In addition, $f^{\prime}(t)=g(t)/(t-1)^{3}$ where $g(t)=2(t-1)-(t+1)\log t$ for $t>0$. One has $g^{\prime}(t)=-(t^{-1}-1-\log(t^{-1}))\leqslant 0$, and since $g(1)=0$, this implies that $g\leqslant 0$ on $(0,1]$ while $g\geqslant 0$ on $1,+\infty)$. Therefore $f^{\prime}(t)=g(t)/(t-1)^{3}\leqslant 0$ for any $t\in\mathbb{R}^{+}\setminus\{1\}$ and thus $f$ is decreasing on $\mathbb{R}^{+}$. We conclude by noting that $\lim_{t\to 0^{+}}f(t)=1$, so that $f\leqslant 1$ on $\mathbb{R}_{+}^{*}$. ∎
@@
For any $p\in\mathbb{R}_{+}^{*}$, the function $q\mapsto D(p,q)$ is strictly convex on $\mathbb{R}_{+}^{*}$, and reaches its minimum (equal to 0) at $q=p$. Hence, it is decreasing on $(0,p]$ and increasing on $p,+\infty)$. In addition, $D(p,q)\leqslant p\log(p/q)$ when $q\leqslant p$, and $D(p,q)\geqslant e^{-1}p\log(p/q)$ when $q\leqslant p/e$.
@@
We may assume, without loss of generality, that $q_{j}>0$ for any $j\in J$. Indeed, either there exists a $j\in J$ such that $q_{j}=0$ and $p_{j}>0$, in which case the left-hand side of (120) is $\infty$ and the inequality holds; or, for any $j\in J$ such that $q_{j}=0$, one has $p_{j}=0$: but in this case replacing $J$ by $J^{\prime}=\{j\in J:q_{j}>0\}$ does not affect the right-hand side of (120).
Next, if $J=\varnothing$, the right-hand side of (120) is 0, thus the inequality holds. We now also assume that $J\neq\varnothing$. Then, by non-negativity and convexity of the function $h$ (Lemma 14), we have
@@
If $q=0$, then by assumption $p=0$ as well, hence all terms in (126) are equal to 0 and the inequalities hold; the same holds when $p=q$. We thus assume $q>0$ and $p\neq q$. In this case, we have $D(p,q)/(\sqrt{p}-\sqrt{q})^{2}=\phi(p/q)$. Now, a direct computation shows that the map $t\mapsto\phi(t^{2})$ is continuously differentiable on $\mathbb{R}_{+}^{*}$, with derivative $t\mapsto 2(t-1)^{-3}\psi(t)$ where $\psi(t)=t^{2}-2t\log t-1$. Since $\psi^{\prime}(t)=2(t-1-\log t)\geqslant 0$ for any $t>0$, the function $\psi$ is non-decreasing, and since $\psi(1)=0$ we deduce that $\psi\leqslant 0$ on $(0,1]$ and $\psi\geqslant 0$ on $1,+\infty)$. Hence $\frac{\mathrm{d}}{\mathrm{d}t}\phi(t^{2})=2(t-1)^{-3}\psi(t)\geqslant 0$ for any $t>0$, $t\neq 1$, thus by continuity of $\phi$ at 0 and 11, the function $\phi$ is non-decreasing on $\mathbb{R}^{+}$.